# Visually Guided Generative Text-Layout Pre-training for Document Intelligence

### v1

@@ Abstract
[…]
@@ 3.2 OCR Evaluation Results
Table 2: OCR text localization and recognition results on SROIE $2019$ benchmark. ${}^{\dagger}$rOCR uses the ground-truth cropped image regions as inputs, whereas ViTLP performs text localization and recognition in a unified stage. All scores are reported in percentage.
@@ 3.3 VDU Evaluation Results
Table 3: VDU evaluation results on form understanding (FUNSD), receipt understanding (CORD), and document classification (RVL-CDIP). ${}^{\dagger}$ FormNet has different sizes of $217$M and $345$M for FUNSD and CORD (Lee et al., 2022). “Maximum Doc-Length” denotes the maximum tokens of an input text sequence that the model can handle.
@@ 3.4.2 Generative Document VQA
Table 5: The results are reported on Average Normalized Levenshtein Similarity (ANLS) between the model-generated answers and ground-truth.
@@ 4 Related Work
LLM-backbone Methods. Most recent studies leverage large language models (LLMs) to tackle multimodal document tasks (Zhang et al., 2024; Ye et al., 2023; Wang et al., 2023). LLaVAR (Zhang et al., 2024) inherits LLaVA architecture (Liu et al., 2023) which directly projects the visual features to LLM embeddings and performs instruction tuning on visual document data. DocLLM (Wang et al., 2023) uses spatial attention to inject $2$D layout information into Llama $2$ (Touvron et al., 2023) with supervised fine-tuning and first enables LLMs to process document information extraction tasks. Thanks to LLMs’ powerful reasoning and generation abilities, utilizing LLMs for visual document processing has become a prominent research trend.
@@ Acknowledgements
We appreciate constructive comments from anonymous ARR reviewers. We thank Bin Liang from CUHK for valuable discussion.
@@ Appendix B Implementation Details of Sequential Layout Head
Figure 6: ViTLP OCR results on a paper. For comprehensive visualization, we render the words and bounding boxes according to ViTLP’s interleaved output sequence. The shown generated OCR results comprise two segments, as the generated tokens reach the decoder sequence length ($M=1024$) in the first segment generation, and the generation process continues by the second segment. Bounding boxes of the first segment are in red, and the second are in green.

### v2

@@ Abstract
Zhiming Mao1,2, Haoli Bai3, Lu Hou3, Jiansheng Wei3, Xin Jiang3, Qun Liu3, Kam-Fai Wong1,2 1The Chinese University of Hong Kong, Hong Kong, China 2MoE Key Laboratory of High Confidence Software Technologies, China 3Noah’s Ark Lab, Huawei Technologies {zmmao, kfwong}@se.cuhk.edu.hk {baihaoli, houlu3, weijiansheng, jiang.xin, qun.liu}@huawei.com
@@ 3.2 OCR Evaluation Results
Table 2: OCR text localization and recognition results on SROIE $2019$ benchmark. TrOCR uses the ground-truth cropped image regions as inputs, whereas ViTLP performs text localization and recognition in a unified stage. All scores are reported in percentage.
@@ 3.3 VDU Evaluation Results
Table 3: VDU evaluation results on form understanding (FUNSD), receipt understanding (CORD), and document classification (RVL-CDIP). FormNet has different sizes of $217$M and $345$M for FUNSD and CORD (Lee et al., 2022). “Maximum Doc-Length” denotes the maximum tokens of an input text sequence that the model can handle.
@@ 3.4.2 Generative Document VQA
Table 5: The results are reported on Average Normalized Levenshtein Similarity (ANLS) between the model-generated answers and ground truth.
@@ 4 Related Work
LLM-backbone Methods. Most recent studies leverage large language models (LLMs) to tackle multimodal document tasks (Zhang et al., 2023; Ye et al., 2023; Wang et al., 2023). LLaVAR (Zhang et al., 2023) inherits LLaVA architecture (Liu et al., 2023) which directly projects the visual features to LLM embeddings and performs instruction tuning on visual document data. DocLLM (Wang et al., 2023) uses spatial attention to inject $2$D layout information into Llama $2$ (Touvron et al., 2023) with supervised fine-tuning and first enables LLMs to process document information extraction tasks. Thanks to LLMs’ powerful reasoning and generation abilities, utilizing LLMs for visual document processing has become a prominent research trend.
@@ Acknowledgements
We appreciate constructive comments from anonymous ARR reviewers. We thank Bin Liang from CUHK for valuable discussion. This research work is partially supported by CUHK direct grant No. $4055209$ and CUHK Knowledge Transfer Project Fund No. KPF23232323GWP20202020.
@@ Appendix B Implementation Details of Sequential Layout Head
Figure 6: ViTLP OCR results on a paper. For comprehensive visualization, we render the words and bounding boxes according to ViTLP’s interleaved output sequence. The shown generated OCR results comprise two segments, as the generated tokens reach the decoder sequence length ($M=1024$) in the first segment generation, and the generation process continues by the second segment. The bounding boxes of the first segment are in red, and the second are in green.